ANALYSIS OF OH^{+}, H_{2}O^{+}, AND H_{3}^{+} IN A DIFFUSE MOLECULAR CLOUD TOWARD W51

\underline{NICK INDRIOLO}, DAVID A. NEUFELD, Department of Physics & Astronomy, Johns Hopkins University, Baltimore, MD 21218; MARYVONNE GERIN, LERMA, CNRS UMR 8112, 24 rue Lhomond, 75231 Paris Cedex 05, France; THOMAS R. GEBALLE, Gemini Observatory, Hilo, HI 96720.

Absorption lines from the molecules OH^{+}, H_{2}O^{+}, and H_{3}^{+} have all been observed in a diffuse molecular cloud along a line of sight near W51 IRS2. We present the first chemical analysis that combines the information provided by all three of these species. Together, OH^{+} and H_{2}O^{+} are used to determine the molecular hydrogen fraction (f_{H_{2}}) in the outskirts of the observed cloud, as well as the product of the cosmic-ray ionization rate of atomic hydrogen and an efficiency factor (ϵζ_{H}). The efficiency factor (ϵ) describes what fraction of the time ionization of H by cosmic rays eventually leads to OH^{+}. H_{3}^{+} is used to infer the cosmic-ray ionization rate of H_{2} (ζ_{2}) in the molecular interior of the cloud. By demanding that the two ionization rates are equal, and taking the value inferred from H_{3}^{+} to be correct, we determine ϵ. This is an important step in the future use of OH^{+} and H_{2}O^{+} on their own as tracers of the cosmic-ray ionization rate.